OpenStudy (anonymous):

find the length of the medians of the triangle with vertices at A(0,0),B(6,0),and C(4,4)

OpenStudy (anonymous):

Find the midpoints of each side. We'll call AB's X (opposite C), BC's Y (opposite A), and AC's Z (opposite B).  Midpoint formula: \[(\frac{x_{1}+x_{2}}{2},\frac{y_{1}+y_{2}}{2})\]

OpenStudy (anonymous):

Distance formula once you find the midpoints: (You'll want distances AY, BZ, and CX) \[d = \sqrt{(x_{2}-x_{1})^2+(y_{2}-y_{1})^2}\] If it helps, the midpoint formula is the average of the coordinates of two points. The distance formula uses the Pythagorean Theorem to measure the 'hypotenuse' (distance) of the right triangle made by two points.